# Improving Adversarial Transferability of Visual-Language Pre-training Models through Collaborative Multimodal Interaction

### v1

@@ II Related Work
The Visual Language Pretraining (VLP) model is a general multimodal understanding model pre-trained on a large-scale dataset of image-text pairs. This model efficiently captures the complex relationships between images and text, establishing deeper and more comprehensive visual and language representations [11]. From the perspective of multimodal fusion, VLP models can be classified into two categories [5]: Fused VLP models and Aligned VLP models. Fused VLP models ($e.g.$, ALBEF [12], TCL [13]) employ a structural approach that initially utilizes a single encoder to extract feature representations from both images and text. Subsequently, a multimodal encoder performs cross-attention operations to achieve feature fusion. It is noteworthy that during the multimodal encoding process, mutual interaction and guidance between image and text features are allowed. Aligned VLP models ($e.g.$, CLIP [14]) utilize a different structural approach. They exclusively use a single encoder to independently learn feature representations for images and text, and subsequently map these features into a shared semantic space. In summary, the VLP model offers a robust foundation for visual and language tasks through its multimodal fusion structure.
@@ IV-B Motivation
By framing the generation of adversarial examples as an optimization problem, the transferability of adversarial examples can be regarded as a manifestation of the model’s generalization ability [26]. Viewing the training of multimodal models, the primary objectives involve achieving modality alignment and constructing a unified visual-language alignment space. Hence, attacks should also consider these aspects. In our analysis, we observe a common issue in current adversarial attacks against VLP models, specifically the lack of corresponding updates in the multimodal feature alignment space during modality updates. Existing attack attempt to expand the distance between images and texts. However, when attacking texts and images, the guiding effect of gradients from images is relatively limited [8], meaning insufficient utilization of cross-modal interaction information. This oversight restricts the capability to capture the subtle changes in such multimodal matching relationships.
@@ IV-C Collaborative Multimodal Interaction Attack
ALBEF PGD $*{}^{*}$ $*{}^{*}$ 3.06 6.79 8.96 13.21 10.34 14.65
BERT-Attack $*{}^{*}$ $*{}^{*}$ 12.64 28.07 29.33 43.17 32.69 46.11
Sep-Attack $*{}^{*}$ $*{}^{*}$ 17.60 32.95 31.17 45.23 32.83 45.49
Co-Attack $*{}^{*}$ $*{}^{*}$ 15.21 29.49 23.60 36.48 25.12 38.89
SGA $*{}^{*}$ $*{}^{*}$ 45.42 55.25 33.38 44.16 34.93 46.57
SGA+MI $*{}^{*}$ $*{}^{*}$ 49.00 58.67 33.99 44.68 37.29 47.65
CMI-Attack (Ours) $*{}^{*}$ $*{}^{*}$ 62.17 69.64 41.84 54.16 46.23 54.68
TCL PGD 6.15 10.78 $*{}^{*}$ $*{}^{*}$ 7.48 13.72 10.34 15.33
BERT-Attack 11.89 26.82 $*{}^{*}$ $*{}^{*}$ 29.69 44.49 33.46 46.06
Sep-Attack 20.13 36.48 $*{}^{*}$ $*{}^{*}$ 31.29 44.65 33.33 45.80
Co-Attack 23.15 40.04 $*{}^{*}$ $*{}^{*}$ 27.85 41.19 30.74 44.11
SGA 48.91 60.34 $*{}^{*}$ $*{}^{*}$ 33.87 44.88 37.74 48.30
SGA+MI 53.60 63.52 $*{}^{*}$ $*{}^{*}$ 35.83 46.10 39.97 49.40
CMI-Attack (Ours) 61.52 71.73 $*{}^{*}$ $*{}^{*}$ 42.45 55.06 48.02 56.09
CLIPViT PGD 2.50 4.93 4.85 8.17 $*{}^{*}$ $*{}^{*}$ 5.36 8.44
BERT-Attack 9.59 22.64 11.80 25.07 $*{}^{*}$ $*{}^{*}$ 30.40 37.43
Sep-Attack 9.59 23.25 11.38 25.60 $*{}^{*}$ $*{}^{*}$ 30.78 39.76
Co-Attack 10.57 24.33 11.94 26.69 $*{}^{*}$ $*{}^{*}$ 32.52 41.82
SGA 13.40 27.22 16.23 30.76 $*{}^{*}$ $*{}^{*}$ 38.76 47.79
SGA+MI 13.87 28.70 16.29 31.02 $*{}^{*}$ $*{}^{*}$ 43.68 49.67
CMI-Attack (Ours) 18.67 37.58 21.18 39.83 $*{}^{*}$ $*{}^{*}$ 49.04 56.71
CLIPCNN PGD 2.09 4.82 4.00 7.81 1.10 6.60 $*{}^{*}$ $*{}^{*}$
BERT-Attack 8.86 23.27 12.33 25.48 27.12 37.44 $*{}^{*}$ $*{}^{*}$
Sep-Attack 8.55 23.41 12.64 26.12 28.34 39.43 $*{}^{*}$ $*{}^{*}$
Co-Attack 8.79 23.74 13.10 26.07 28.79 40.03 $*{}^{*}$ $*{}^{*}$
SGA 11.42 24.80 14.91 28.82 31.24 42.12 $*{}^{*}$ $*{}^{*}$
SGA+MI 12.30 25.40 15.31 29.58 31.29 42.69 $*{}^{*}$ $*{}^{*}$
CMI-Attack (Ours) 18.87 37.98 22.23 40.07 39.39 52.29 $*{}^{*}$ $*{}^{*}$
TABLE I: Attack success rate (%) in image-text retrieval on the Flickr30K dataset. We present the attack success rate metric R@1 for both IR and TR tasks, indicating the success of attacks at Rank 1. $*{}^{*}$ indicates white-box attacks.
@@
ALBEF PGD $*{}^{*}$ $*{}^{*}$ 12.46 17.77 13.96 23.10 17.45 23.54
BERT-Attack $*{}^{*}$ $*{}^{*}$ 24.34 33.39 44.94 52.28 47.73 54.75
Sep-Attack $*{}^{*}$ $*{}^{*}$ 32.83 42.92 44.03 54.46 46.96 55.88
Co-Attack $*{}^{*}$ $*{}^{*}$ 32.62 43.09 44.89 54.75 47.30 55.64
SGA $*{}^{*}$ $*{}^{*}$ 58.56 65.38 57.06 65.25 58.95 66.52
SGA+MI $*{}^{*}$ $*{}^{*}$ 60.82 67.02 59.69 65.88 59.79 67.54
CMI-Attack (Ours) $*{}^{*}$ $*{}^{*}$ 72.09 75.57 64.56 71.12 66.37 72.59
TCL PGD 10.83 16.52 $*{}^{*}$ $*{}^{*}$ 14.23 22.28 17.25 23.12
BERT-Attack 35.32 45.92 $*{}^{*}$ $*{}^{*}$ 51.09 58.80 52.23 61.26
Sep-Attack 41.71 52.97 $*{}^{*}$ $*{}^{*}$ 50.74 60.13 51.90 61.26
Co-Attack 46.08 57.09 $*{}^{*}$ $*{}^{*}$ 51.62 60.46 52.13 62.49
SGA 65.93 73.30 $*{}^{*}$ $*{}^{*}$ 56.34 63.99 59.44 65.70
SGA+MI 68.68 76.08 $*{}^{*}$ $*{}^{*}$ 57.69 65.00 59.99 67.12
CMI-Attack (Ours) 78.63 82.55 $*{}^{*}$ $*{}^{*}$ 66.65 71.91 68.66 74.01
CLIPViT PGD 7.24 10.75 10.19 13.74 $*{}^{*}$ $*{}^{*}$ 7.32 11.34
BERT-Attack 20.34 29.74 21.08 29.61 $*{}^{*}$ $*{}^{*}$ 44.54 55.32
Sep-Attack 23.41 34.61 25.77 36.84 $*{}^{*}$ $*{}^{*}$ 43.11 49.76
Co-Attack 30.28 42.67 32.84 44.69 $*{}^{*}$ $*{}^{*}$ 55.08 62.51
SGA 33.41 44.64 37.54 47.76 $*{}^{*}$ $*{}^{*}$ 58.93 65.83
SGA+MI 34.93 45.70 35.87 46.68 $*{}^{*}$ $*{}^{*}$ 61.26 67.84
CMI-Attack (Ours) 48.41 58.26 49.10 60.16 $*{}^{*}$ $*{}^{*}$ 72.62 76.07
CLIPCNN PGD 7.01 10.62 10.08 13.65 4.88 10.70 $*{}^{*}$ $*{}^{*}$
BERT-Attack 23.38 34.64 24.58 29.61 51.28 57.49 $*{}^{*}$ $*{}^{*}$
Sep-Attack 26.53 39.29 30.26 41.51 50.44 57.11 $*{}^{*}$ $*{}^{*}$
Co-Attack 29.83 41.97 32.97 43.72 53.10 58.90 $*{}^{*}$ $*{}^{*}$
SGA 31.61 43.00 34.81 45.95 56.62 60.77 $*{}^{*}$ $*{}^{*}$
SGA+MI 31.84 43.38 33.23 44.77 57.99 60.98 $*{}^{*}$ $*{}^{*}$
CMI-Attack (Ours) 45.30 55.62 47.78 57.94 67.72 70.83 $*{}^{*}$ $*{}^{*}$
TABLE II: Attack success rate (%) in image-text retrieval on the MSCOCO Dataset. We present the attack success rate metric R@1 for both IR and TR tasks, indicating the success of attacks at Rank 1. $*{}^{*}$ indicates white-box attacks.
@@ V-A Setup
Models. We study two main types of visual language models: fused VLP models and aligned VLP models. The fused models early fused visual and language features when processing two modal data, and achieved joint processing through shared layers. ALBEF [12] and TCL [13] are representative fused models, both of which use 12-layer ViT-B/16 [27] as the image encoder, and two independent 6-layer Transformers for text and multimodal encoding. The main difference is their distinct pre-training objectives. On the contrary, the aligned model initially independently processed images and text, and then aligned multimodal representations in the deep layer to learn the complex relationship between the two. We also examine two variants of CLIP [14], namely $\text{CLIP}_{\text{ViT}}$ and $\text{CLIP}_{\text{CNN}}$. The former is based on ViT-B/16 image encoder, while the latter uses ResNet-101 [28] as the backbone network. By evaluating the transferability of adversarial examples between the source and the target model, we verify the effectiveness of the proposed framework.

### v2

# Improving Adversarial Transferability of Vision-Language Pre-training Models through Collaborative Multimodal Interaction

@@ II Related Work
The Vision-Language Pretraining (VLP) model is a general multimodal understanding model pre-trained on a large-scale dataset of image-text pairs. This model efficiently captures the complex relationships between images and text, establishing deeper and more comprehensive visual and language representations [11]. From the perspective of multimodal fusion, VLP models can be classified into two categories [5]: Fused VLP models and Aligned VLP models. Fused VLP models ($e.g.$, ALBEF [12], TCL [13]) employ a structural approach that initially utilizes a single encoder to extract feature representations from both images and text. Subsequently, a multimodal encoder performs cross-attention operations to achieve feature fusion. It is noteworthy that during the multimodal encoding process, mutual interaction and guidance between image and text features are allowed. Aligned VLP models ($e.g.$, CLIP [14]) utilize a different structural approach. They exclusively use a single encoder to independently learn feature representations for images and text, and subsequently map these features into a shared semantic space. In summary, the VLP model offers a robust foundation for visual and language tasks through its multimodal fusion structure.
@@ IV-B Motivation
By framing the generation of adversarial examples as an optimization problem, the transferability of adversarial examples can be regarded as a manifestation of the model’s generalization ability [26]. Viewing the training of multimodal models, the primary objectives involve achieving modality alignment and constructing a unified vision-language alignment space. Hence, attacks should also consider these aspects. In our analysis, we observe a common issue in current adversarial attacks against VLP models, specifically the lack of corresponding updates in the multimodal feature alignment space during modality updates. Existing attack attempt to expand the distance between images and texts. However, when attacking texts and images, the guiding effect of gradients from images is relatively limited [8], meaning insufficient utilization of cross-modal interaction information. This oversight restricts the capability to capture the subtle changes in such multimodal matching relationships.
@@ IV-C Collaborative Multimodal Interaction Attack
ALBEF PGD 52.45 58.65 3.06 6.79 8.96 13.21 10.34 14.65
BERT-Attack 11.57 27.46 12.64 28.07 29.33 43.17 32.69 46.11
Sep-Attack 65.69 73.95 17.60 32.95 31.17 45.23 32.83 45.49
Co-Attack 77.16 83.86 15.21 29.49 23.60 36.48 25.12 38.89
SGA 97.24 97.28 45.42 55.25 33.38 44.16 34.93 46.57
SGA+MI 96.14 96.40 49.00 58.67 33.99 44.68 37.29 47.65
CMI-Attack (Ours) 97.08 97.43 62.17 69.64 41.84 54.16 46.23 54.68
TCL PGD 6.15 10.78 77.87 79.48 7.48 13.72 10.34 15.33
BERT-Attack 11.89 26.82 14.54 29.17 29.69 44.49 33.46 46.06
Sep-Attack 20.13 36.48 84.72 86.07 31.29 44.65 33.33 45.80
Co-Attack 23.15 40.04 77.94 85.59 27.85 41.19 30.74 44.11
SGA 48.91 60.34 98.37 98.81 33.87 44.88 37.74 48.30
SGA+MI 53.60 63.52 97.26 97.98 35.83 46.10 39.97 49.40
CMI-Attack (Ours) 61.52 71.73 98.00 98.67 42.45 55.06 48.02 56.09
CLIPViT PGD 2.50 4.93 4.85 8.17 70.92 78.61 5.36 8.44
BERT-Attack 9.59 22.64 11.80 25.07 28.34 39.08 30.40 37.43
Sep-Attack 9.59 23.25 11.38 25.60 79.75 86.79 30.78 39.76
Co-Attack 10.57 24.33 11.94 26.69 93.25 95.86 32.52 41.82
SGA 13.40 27.22 16.23 30.76 99.08 98.94 38.76 47.79
SGA+MI 13.87 28.70 16.29 31.02 99.26 99.10 43.68 49.67
CMI-Attack (Ours) 18.67 37.58 21.18 39.83 98.16 98.81 49.04 56.71
CLIPCNN PGD 2.09 4.82 4.00 7.81 1.10 6.60 86.46 92.25
BERT-Attack 8.86 23.27 12.33 25.48 27.12 37.44 30.40 40.10
Sep-Attack 8.55 23.41 12.64 26.12 28.34 39.43 91.44 95.44
Co-Attack 8.79 23.74 13.10 26.07 28.79 40.03 94.76 96.89
SGA 11.42 24.80 14.91 28.82 31.24 42.12 99.24 99.49
SGA+MI 12.30 25.40 15.31 29.58 31.29 42.69 99.62 99.74
CMI-Attack (Ours) 18.87 37.98 22.23 40.07 39.39 52.29 99.74 99.76
TABLE I: Attack success rate (%) in image-text retrieval on the Flickr30K dataset. We present the attack success rate metric R@1 for both IR and TR tasks, indicating the success of attacks at Rank 1. indicates white-box attacks.
@@
ALBEF PGD 76.70 86.30 12.46 17.77 13.96 23.10 17.45 23.54
BERT-Attack 24.39 36.13 24.34 33.39 44.94 52.28 47.73 54.75
Sep-Attack 82.60 89.88 32.83 42.92 44.03 54.46 46.96 55.88
Co-Attack 79.87 87.83 32.62 43.09 44.89 54.75 47.30 55.64
SGA 96.75 96.95 58.56 65.38 57.06 65.25 58.95 66.52
SGA+MI 96.03 96.26 60.82 67.02 59.69 65.88 59.79 67.54
CMI-Attack (Ours) 97.40 97.51 72.09 75.57 64.56 71.12 66.37 72.59
TCL PGD 10.83 16.52 59.58 69.53 14.23 22.28 17.25 23.12
BERT-Attack 35.32 45.92 38.54 48.48 51.09 58.80 52.23 61.26
Sep-Attack 41.71 52.97 70.32 78.97 50.74 60.13 51.90 61.26
Co-Attack 46.08 57.09 85.38 91.39 51.62 60.46 52.13 62.49
SGA 65.93 73.30 98.97 99.15 56.34 63.99 59.44 65.70
SGA+MI 68.68 76.08 98.89 98.87 57.69 65.00 59.99 67.12
CMI-Attack (Ours) 78.63 82.55 98.94 99.30 66.65 71.91 68.66 74.01
CLIPViT PGD 7.24 10.75 10.19 13.74 54.79 66.85 7.32 11.34
BERT-Attack 20.34 29.74 21.08 29.61 45.06 51.68 44.54 55.32
Sep-Attack 23.41 34.61 25.77 36.84 68.52 77.94 43.11 49.76
Co-Attack 30.28 42.67 32.84 44.69 97.98 98.80 55.08 62.51
SGA 33.41 44.64 37.54 47.76 99.79 99.79 58.93 65.83
SGA+MI 34.93 45.70 35.87 46.68 99.77 98.67 61.26 67.84
CMI-Attack (Ours) 48.41 58.26 49.10 60.16 99.47 99.73 72.62 76.07
CLIPCNN PGD 7.01 10.62 10.08 13.65 4.88 10.70 76.99 84.20
BERT-Attack 23.38 34.64 24.58 29.61 51.28 57.49 54.43 62.17
Sep-Attack 26.53 39.29 30.26 41.51 50.44 57.11 88.72 92.49
Co-Attack 29.83 41.97 32.97 43.72 53.10 58.90 96.72 98.56
SGA 31.61 43.00 34.81 45.95 56.62 60.77 99.61 99.80
SGA+MI 31.84 43.38 33.23 44.77 57.99 60.98 99.71 99.92
CMI-Attack (Ours) 45.30 55.62 47.78 57.94 67.72 70.83 99.75 99.95
TABLE II: Attack success rate (%) in image-text retrieval on the MSCOCO Dataset. We present the attack success rate metric R@1 for both IR and TR tasks, indicating the success of attacks at Rank 1. indicates white-box attacks.
@@ V-A Setup
Models. We study two main types of vision-language models: fused VLP models and aligned VLP models. The fused models early fused visual and language features when processing two modal data, and achieved joint processing through shared layers. ALBEF [12] and TCL [13] are representative fused models, both of which use 12-layer ViT-B/16 [27] as the image encoder, and two independent 6-layer Transformers for text and multimodal encoding. The main difference is their distinct pre-training objectives. On the contrary, the aligned model initially independently processed images and text, and then aligned multimodal representations in the deep layer to learn the complex relationship between the two. We also examine two variants of CLIP [14], namely $\text{CLIP}_{\text{ViT}}$ and $\text{CLIP}_{\text{CNN}}$. The former is based on ViT-B/16 image encoder, while the latter uses ResNet-101 [28] as the backbone network. By evaluating the transferability of adversarial examples between the source and the target model, we verify the effectiveness of the proposed framework.